LASER INDUCED FLUORESCENCE SPECTROSCOPY OF THE KETENYL RADICAL

L. R. BROCK and \underline{ERIC A. ROHLFING}^{a}, Combustion Research Facility, Sandia National Laboratories, Livermore, CA; B. MISCHLER, Paul Scherrer Institut, Ch-5232, Viligen PSI, Switzerland.

The ketenyl radical, HCCO, is widely recognized as a critical reaction intermediate in hydrocarbon combustion. In order to enable laser-based diagnostics of ketenyl in combustion and kinetics environments, we have examined the electronic spectroscopy of the B ^{2}Π - X ^{2}A” band system of H(D)CCO using laser-induced fluorescence in a free-jet environment. We obtain both vibronically resolved excitation spectra and fully rotationally resolved spectra for the origin band and for four bands terminating on levels of Σ vibronic symmetry, which are assigned to the two pairs of Σ states derived from one quantum of excitation in each of the Renner-Teller active modes, the CCO and CCH(D) bend. In addition to the Renner-Teller and spin-orbit couplings, there is substantial evidence for additional perturbations among the low-lying bending levels in the first excited state of ketenyl.

aThis work supported by the U.S. Department of Energy, Office of Basic Energy Sciences, Division of Chemical Sciences